OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. ? inches Part B How many miles are represented by 1/4 inch on the map? Enter your answer in the box. ? miles

OpenStudy (anonymous):

please help? I think I lost my self? :/

OpenStudy (anonymous):

Set up a proportion: 2/25=x/60 Multiply 60 to both sides: 60×2/25 Simplify

OpenStudy (anonymous):

of course 4

OpenStudy (anonymous):

I think that is what it should be.

OpenStudy (anonymous):

but the second one for B ? is it 3? or 3.125

OpenStudy (anonymous):

thats the one im really confused on :/ :(

OpenStudy (anonymous):

For B, divide 25 by 2, then divide by 1/4 I think

OpenStudy (anonymous):

its 3.125?

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

I have to look at my paper

OpenStudy (anonymous):

okay so you know that 2inch=25miles you want a quarter inch what fraction is a quarter equal to, where you have 2 as the numerator 2/8 is the same as 1/4 so if you know 2inches=25 miles then 2/8 is the same as 25/8 so your answer is 25/8 miles